OpenStudy (vera_ewing):

Solve the following system of equations: 2x + 3y – z = 1 3x + y + 2z = 12 x + 2y – 3z = –5 (5 points) (3, 1, 2) (–3, 1, 2) (3, –1, 2) (3, 1, –2)

OpenStudy (vera_ewing):

@wio I got A am I right?

OpenStudy (jdoe0001):

heheh

OpenStudy (jdoe0001):

makes one wonder how A came about

OpenStudy (vera_ewing):

Well I just plugged in the points and they worked for A so I want to see what others got to make sure I'm right

OpenStudy (amistre64):

if A makes them all true, then A it is :)

OpenStudy (jdoe0001):

hmmm I guess that's one way to go about it

OpenStudy (michele_laino):

hint: try to substitute the values in A into your system

OpenStudy (vera_ewing):

Did you get A as well?

OpenStudy (jdoe0001):

just bear in mind, the choices are there as guidelines, not as a menu to pick from chicken or beef or turkey or whatever else

OpenStudy (amistre64):

rref{{2,3,-1,1},{3,1,2,12},{1,2,-3,-5}} im not getting A and yes, trial and error is a valid method.

OpenStudy (vera_ewing):

the points can be plugged in for x, y, and z respectively? and are you sure A's points don't work?

OpenStudy (amistre64):

show your work, plug them in and ill show you your error

OpenStudy (vera_ewing):

2(3) + 3(-1) – (2) = 1 6-3-2=1 3-2=1 1=1 3x + y + 2z = 12 x + 2y – 3z = –5 (3, –1, 2) (3, 1, –2)

OpenStudy (amistre64):

so A is not y=-1

OpenStudy (vera_ewing):

right, so it's not A or B. So is it C or D?

OpenStudy (amistre64):

i agree its C or D

OpenStudy (vera_ewing):

how do i figure out if it's C or D?

OpenStudy (amistre64):

well, the same way your approach is working, plug in the values and chk

OpenStudy (amistre64):

plug in A, it fails plug in B it fails plug in C ... ??? 2(3) + 3(-1) – (2) = 1 3(3) + (-1) + 2(2) = 12 (3) + 2(-1) – 3(2) = –5 6-3-2 = 1 9-1+4 = 12 3-2-6 = –5 are these true?

OpenStudy (vera_ewing):

yes, the values for C are true.

OpenStudy (amistre64):

umm, no. they have to equal what they equal .... their equality is already defined by the equations that exist in the system. then since C fits, C must be the solution.

OpenStudy (vera_ewing):

yeah sorry i made a mistake haha thank you

OpenStudy (amistre64):

good luck :)